OpenStudy (anonymous):

How do you graph 3 to the power of x?

OpenStudy (nory):

Here is kind of the general shape of an exponential function: 

OpenStudy (anonymous):

Ok, but if you were to make a table of points, what would some of those points be?

OpenStudy (nory):

Well, f(x) = 3^x. As you know from basic algebra, 3^0 = 1, and 3^1 = 3. And 3^2 = 9...and so on.

OpenStudy (nory):

So you have some points: (0,1) , (1,3) , (2,9). By following this same procedure, you can find more easily. You just raise 3 to different powers and write down the results.

OpenStudy (anonymous):

ok, thanks

OpenStudy (nory):

It is interesting because the function starts off pretty slow...1, 3, 9... but then really takes off. 27, 81, 243...

OpenStudy (nory):

(Just an interesting tidbit.)

OpenStudy (anonymous):

So I really just need like three points and then just have the line go up for a while